# Homomorphism, substructure and ideal: Elementary but rigorous aspects of renormalization group or hierarchical structure of topological orders

### v1

@@ II Substructure from homomorphism and projection as renormalization group flow
Figure 3: Hierarchical structure of gapped phase induced from massless RG. The homomorphism $\rho$ induces natural hierarchical structures in $1+1$ dimensional gapped phase and corresponding $2+1$ dimensional bulk states of TOs. Assuming Moore-Seiberg data, one can label primary states by the object in $\mathbf{A}_{(1)}$ or $\mathbf{A}_{(2)}$. However, the label of the states in BCFT can be outside of $\mathbf{A}_{(1)}$ and called symmetry breaking boundary states or new boundary states and they will be labeled by extended algebra $\mathbf{A}_{(1)}^{\text{ex}}$ or $\mathbf{A}_{(2)}^{\text{ex}}$.
@@
When the system $(1)$ preserves the symmetry $\mathbf{A}^{\prime}_{(1)}\subset\mathbf{A}_{(1)}$ one can investigate possible Hilbert space as $\mathbf{A^{\prime}}_{(1)}$ invariant Hilbert space or module (not states) spanned by these BCFTs. The spontaneous $\mathbf{A}^{\prime}_{(1)}$ symmetry breaking corresponds to the appearance of $\mathbf{A}^{\prime}_{(1)}$ noninvariant states. The existence of massless flow $\rho:\mathbf{A}_{(1)}\rightarrow\mathbf{A}_{(2)}$ implies the hierarchical relation of the symmetry of gapped phase as $\rho:\mathbf{A}^{\prime}_{(1)}\rightarrow\mathbf{A}^{\prime}_{(2)}$ and that of the corresponding modules. By applying the argument of the construction of projection or embedding $\mathbf{A}_{(2)}\subset\mathbf{A}_{(1)}$, one can also obtain the corresponding projection for $\mathbf{A^{\prime}}_{(2)}\subset\mathbf{A^{\prime}}_{(1)}$ and corresponding hierarchy of gapped Hamiltonian. In other words, the massless RG almost automatically implies a hierarchical structure of $1+1$ dimensional gapped phases. Moreover, this provides hierarchical structures of entanglement spectrum or bulk states of $2+1$ dimensional topologically ordered systems by considering the CFT/TQFT correspondence[39, 163, 164, 165, 166, 167, 168, 22] (FIG. 3).
@@
where $\beta_{(1)}$ is an element of the ring $\mathbf{A}_{(1)}^{\text{ex}}$. By restricting $\beta_{(1)}\in\mathbf{A^{\prime}}_{(1)}$ and choosing modules labeled by the elements of extended algebra $\mathbf{A}_{(1)}^{\text{ex}}$, one can obtain appropriate $\mathbf{A}^{\prime}_{(2)}$ module labeled by the elements of the extended algebra $\mathbf{A}_{(2)}^{\text{ex}}$. This argument provides a symmetry-based formalism to characterize the hierarchical structure of the gapped phases revealing the symmetry properties of Hilbert spaces. For further studies in this research direction, the construction of symmetry-breaking boundary conditions and the corresponding defects are fundamentally important.
@@ II.1 Unbroken and emergent symmetry
Based on the arguments in this section, we provide a definition of unbroken (or exact) symmetry and emergent symmetry under an RG flow $\rho:\mathbf{A}_{(1)}\rightarrow\mathbf{A}_{(2)}$. A phenomenology of unbroken symmetry under RG flow is simple that a symmetry subring $\mathbf{A^{\prime}}_{(1)}\subset\mathbf{A}_{(1)}$ of UV theory which does not change under a RG flow $\rho$ is an unbroken symmetry. In the algebraic language, this condition means that $\mathbf{A^{\prime}}_{(1)}$ and $\text{Ker}\rho$ are disjoint except for 00. Hence the following defines the unbroken symmetry,
@@
$\begin{split}&\mathbf{A^{\prime}}_{(1)}\text{ is a unbroken symmetry under }\rho\\ &\leftrightarrow\mathbf{A^{\prime}}_{(1)}\cap\text{Ker}\rho=\{0\}\end{split}$ (7)
@@
In other words, when the above condition is not satisfied, the symmetry $\mathbf{A^{\prime}}_{(1)}$ is deformed or broken nontrivially under $\rho$. This general argument provides a testable and accessible criterion of emergent and exact symmetry and will be useful to test conjectural RG flow, for example in [121, 122, 123]. It should be remarked that as linear algebra, one can consider the basis of $\mathbf{A}_{(1)}/\text{Ker}\rho$, but it does not imply that $\mathbf{A}_{(1)}/\text{Ker}\rho$ forms subring of original ring $\mathbf{A}_{(1)}$. Hence the dimension of unbroken symmetry (which should form subalgebra) is less than the dimension of $\mathbf{A}_{(1)}/\text{Ker}\rho$ and the following inequality holds,
$\text{dim}\mathbf{A}_{(2)}\geq\text{dim}\mathbf{A^{\prime}}_{(1)},$ (8)
@@
where dim is the dimension of a ring as a linear algebra. By taking a maximal unbroken subalgebra under $\rho$, one can restrict the form of perturbation in the UV theory and the above inequality provides the constraint in comparing the IR (or emergent) theory.
@@
For the readers interested in the perturbative quantum field theoretic description of our method, we note two possible research directions. The first is a straightforward analysis of the interpolation behavior of $H_{(1)}$ to $H_{(2)}$. This interpolation can be implemented by preparing the Hamiltonian $(1-\lambda)H_{(1)}+\lambda H_{(2)}$ with $\lambda=0$ to $\lambda=1$ and the perturbative analysis around $\lambda=0$ is useful also in QFTs. In this research direction, numerical methods such as truncated conformal space approach[182, 183] and analysis of corresponding lattice models will be fundamentally important.
@@
The second is the detection of maximal unbroken subalgebra in this subsection. By assuming that a general perturbation $H_{\text{per}}$ from a UV CFT (possibly, multiple perturbations in QFTs) is invariant under the symmetry corresponding to this subalgebra $\mathbf{A^{\prime}}_{(1)}$, one can restrict the form of possible perturbations systematically. More conventionally, one can write down the condition as follows,
$[H_{\text{per}},Q_{a^{\prime}_{(1)}}]=0,$ (9)
@@
where $Q_{a^{\prime}_{(1)}}$ is a symmetry operator (or integral of motions) corresponding to the element $a^{\prime}_{(1)}\in\mathbf{A^{\prime}}_{(1)}$. Inversely, if the perturbation $H_{\text{per}}$ is given, one can determine possible RG flows by studying the maximal subalgebra structure which commutes with $H_{\text{per}}$ and its disjoint ideals. Related arguments based on group symmetry can be seen ubiquitously, and our argument can be considered as a generalization emphasizing the ideal structure. We note [184, 185] as references containing analysis of symmetry with ring structure and their action to bulk fields.
@@ II.2 Historical remarks on the projections in physics
Before moving to the next section, we comment on a history of the ideas to represent RG as a projection. The idea itself can be seen ubiquitously in the fields, but well-known earlier examples related to the discussions in this work are the corner-transfer method by Baxter [186](for more details, see the review[187]) and projective representation in Affleck-Kenedy-Lieb-Tasaki chain[47]. This kind of projection or projective representation played a significant role in studying many-body quantum systems both theoretically and numerically. Related historical aspects can be seen in the review [188].
Related to the argument in CFT in this manuscript, we also note the restriction of Hilbert space in quantum group symmetric models[189, 190, 191] which have a connection to the Dotsenko-Fateev Coulomb gas[192, 193] (or Fegin-Fuchs construction[194] based on Wakimoto free field representation[195]) and the massless RG flow[87]. More recently, the RG domain wall, which is in the main scope of this work has been introduced in [32] and its general correspondence with existing coset CFTs has been proposed in [33]. Some of the properties of homomorphism, or property of symmetry-preserving domain wall, have already appeared in [33] and it is remarkable that this pioneering work appeared before the introduction of generalized symmetry[196]. Later, more general cases have been studied in [197, 198, 199, 200, 201] and more mathematical understanding has been developed in [118, 119, 113, 202, 68, 203, 204, 205]. The idea of RG domain wall has also been combined with the holography in [206, 207] and related bound of entropy of conformal interface has been discussed in [208]. More recently, the correspondence between symmetry-preserving domain wall[23, 24, 25, 26, 28] and RG domain wall has been realized in [203, 67, 88, 209, 69], but this is relatively less familiar in the fields.
@@ III Generating ideal: Role of nonabelian anyon or noninvertible symmetry
One can perform the exact construction by classifying ideals of a UV theory and obtain (infinitely) many new series of unexplored homomorphism. In other words, if the above statement does not hold, the connection between the RG domain wall and massless flow will be broken. Further investigation of more detailed properties of the IR theory, such as conditions for the locality or nonlocality and conditions to obtain IR theory with positive or integer fusion coefficients, is an open problem. However, surprisingly, we note that the related problems of the projection for Yang-Lee CFT flows have already been investigated much earlier in [222]. The categorical analog of the ideal called additive congruence relation by ideal will play a role in the more categorical description of our method [^1] We thank Jurgen Fuchs for teaching the terminology to us.. Related to this research direction, we remark that the idea of ideal appears in the studies of Hopf algebra and related TQFTs (see [224, 225], for example). We also remark that in the studies on Hopf algebra, one can see the ideas to represent a RG as projection [226, 227, 228, 229] but their formalism is closer to Lagrangian formalism and its precise connection to RG domain wall is (reasonable but) unclear at this stage (see also the reviews[230, 231])
Figure 4: Picture of the noninvertible object (anyon or symmetry) as a seed to generate a condensable block of a theory. We used blue color for noninvertible objects black color for fusion between the noninvertible object and other objects and red color for objects in $\mathbf{A}_{1}$. Because of the noninvertible structure, one can send the ideal generated by $s_{(1)}$ to 00 only by applying a projection to the theory.
@@
In particular, if one takes a noninvertible object in $\mathbf{A}_{(1)}$, it generates a nontrivial ideal straightforwardly because the generated ideal cannot contain identity. From this view, one can say that a noninvertible symmetry corresponding to nonabelian anyon is a seed to obtain ideal and corresponding homomorphism or massless RG flow (FIG. 4). By combining this with Moore-Seiberg data[163, 164, 165, 166], the fundamental importance of further systematic studies on noninvertible symmetries (or integral of motions) and the corresponding nonabelian anyons for the general understanding of TOs becomes more evident.
One can construct possible IR theories only from the algebraic data of a UV theory, but the construction of the homomorphisms between given UV and IR theory requires more involved calculation and consistency check. In the next section, we solve this problem in several models.
@@ IV Construction of homomorphism
In this section, we study ring homomorphisms between fusion rings appearing massless RG flow of conformal field theories[87]. We study the massless RG flow between ultraviolet (UV) CFT $\mathbf{B}$ to infrared (IR) CFT $\mathbf{B}^{\prime}$ where $\mathbf{B}$ and $\mathbf{B}^{\prime}$ represents nonchiral fusion ring or anyon algebra of spherical fusion category (SFC)[232, 233, 155]. Roughly speaking, the SFC corresponds to the bosonic modular invariant and conformal bootstrap technique[234]. We concentrate our attention on the algebraic relation of objects or anyons, study the ring homomorphism $\rho$: $\mathbf{B}$rightarrow $\mathbf{B}^{\prime}$ (or algebraic data of functor between two SFCs[22, 24, 26]). To simplify the notation and discussion, we assume the theory is in the scope of Moore-Seiberg data[163, 164, 165, 166] and the ring homomorphism between nonchiral fusion rule naturally introduces ring homomorphism between chiral fusion rule and vise-versa. Algebraically, this is trivial because of the ring isomorphism between the nonchiral and chiral fusion rule.
Before moving into the respective discussions, we remark on one of the most nontrivial points of our findings. Surprisingly, even in the flow from the tricritical Ising model to the Ising model, the homomorphism is not unique. However, by requiring a variant of integrability, the solution becomes unique. This phenomenon itself is analogous to ideal number decomposition by Kummer and the integrability will play a fundamental role in constructing a treatable class of models. Inversely, the other homomorphisms are mysterious to some extent but there exists a conjectural phenomenology to them. If one assumes the IR theory $\mathbf{A}_{(2)}$ or consequence of RG flow is integrable, one can say this homomorphism induces partially integrable embedding of the theory to UV theory $\mathbf{A}_{(1)}$, $\mathbf{A}_{(2)}\subset\mathbf{A}_{(1)}$. Fortunately, there exist very similar phenomena called partially solvable systems in recent literature[92, 93, 94] [^2] We thank Chihiro Matsui for introducing us to the idea of this embedding and related ideas in the literature and Hosho Katsura for indicating useful references in this research field.. Historical aspects of those models have been summarized concisely in the introduction of [94], and we note the Hilbert space fragmentation as a fundamental aspect and note a recent review [236]. We conjecture that ring homomorphism generally corresponds to integrable or partially solvable systems. Further analysis of the RG flows by perturbative QFT methods is an interesting direction and it will be described by UV CFT with multiple perturbations as in [237].

### v2

@@ II Substructure from homomorphism: Projection as renormalization group flow
Figure 3: Hierarchical structure of gapped phase induced from massless RG. The homomorphism $\rho$ induces natural hierarchical structures in $1+1$ dimensional gapped phase and corresponding $2+1$ dimensional bulk states of TOs. Assuming Moore-Seiberg data, one can label primary states by the object in $\mathbf{A}_{(1)}$ or $\mathbf{A}_{(2)}$. However, the label of the states in BCFT can be outside of $\mathbf{A}_{(1)}$ and called symmetry breaking boundary states or new boundary states and they will be labeled by extended algebra $\mathbf{A}_{(1)}^{\text{ex}}$ or $\mathbf{A}_{(2)}^{\text{ex}}$. Without such extensions, they will be labelled by objects in an ideal of $\mathbf{A^{\prime}}_{(1)}$ or $\mathbf{A^{\prime}}_{(2)}$.
@@
When the system $(1)$ preserves the symmetry $\mathbf{A}^{\prime}_{(1)}\subset\mathbf{A}_{(1)}^{\text{ex}}$ one can investigate possible Hilbert space as $\mathbf{A^{\prime}}_{(1)}$ invariant Hilbert space or module (not states) spanned by these BCFTs. The spontaneous $\mathbf{A}^{\prime}_{(1)}$ symmetry breaking corresponds to the appearance of $\mathbf{A}^{\prime}_{(1)}$ noninvariant states. The existence of massless flow $\rho:\mathbf{A}_{(1)}\rightarrow\mathbf{A}_{(2)}$ implies the hierarchical relation of the symmetry of the gapped phase as $\rho:\mathbf{A}^{\prime}_{(1)}\rightarrow\mathbf{A}^{\prime}_{(2)}$ and that of the corresponding modules. By applying the argument of the construction of projection or embedding $\mathbf{A}_{(2)}\subset\mathbf{A}_{(1)}$, one can also obtain the corresponding projection for $\mathbf{A^{\prime}}_{(2)}\subset\mathbf{A^{\prime}}_{(1)}$ and corresponding hierarchy of gapped Hamiltonian. In other words, the massless RG almost automatically implies a hierarchical structure of $1+1$ dimensional gapped phases. Moreover, this provides hierarchical structures of entanglement spectrum or bulk states of $2+1$ dimensional topologically ordered systems by considering the CFT/TQFT correspondence[38, 165, 166, 167, 168, 169, 170, 21] (FIG. 3).
@@
where $\beta_{(1)}$ is an element of the ring $\mathbf{A}_{(1)}^{\text{ex}}$. By restricting $\beta_{(1)}\in\mathbf{A^{\prime}}_{(1)}$ and choosing modules labeled by the elements of extended algebra $\mathbf{A}_{(1)}^{\text{ex}}$, one can obtain appropriate $\mathbf{A}^{\prime}_{(2)}$ module labeled by the elements of the extended algebra $\mathbf{A}_{(2)}^{\text{ex}}$. This argument provides a symmetry-based formalism to characterize the hierarchical structure of the gapped phases, revealing the symmetry properties of Hilbert spaces. For further studies in this research direction, the construction of symmetry-breaking boundary conditions and the corresponding defects is fundamentally important.
We also note that if one assumes that the BCFTs are labelled by $\mathbf{A^{\prime}}$, one can observe that a $\mathbf{A^{\prime}}$ module is labelled by objects in an ideal of $\mathbf{A^{\prime}}$. For the simplest example, one can obtain the trivial ideal $\mathbf{A^{\prime}}$. In other words,
A gapped phase is an ideal of unbroken subring A’. (8)
when assuming this modified version of the Moore-Seiberg data and massive RG flow. Inversely, $\mathbf{A^{\prime}}$ invariant sets $\mathbf{R}$ in the extended ring $\mathbf{A}^{\text{ex}}$ can be thought of as a generalization of an ideal.
Based on the recent development of sandwich construction, one can interpret the set $\mathbf{R}$ as follows, in general,
$\begin{split}&\{\text{Label of $D$-dimensional gapped phase}\}\\ =&\{\text{Condensable $D+1$-dimension symmetry}\}\end{split}$ (9)
where both $D$ and $D+1$ dimensional system has same symmetry described by a ring $\mathbf{A}^{\prime}$. The latter also determines the massless RG flows in $D$-dimensional systems with the UV symmetry $\mathbf{A}^{\prime}$. In other words, the condensable objects in a $D+1$-dimensional system can appear in a $D$-dimensional boundary as stable (or surviving) objects, and this phenomenon is governed by an ideal or its generalizations. This provides a more complete and general algebraic understanding of the sandwich construction proposed in [90].
@@ II.1 Unbroken and emergent symmetry
Based on the arguments in this section, we provide a definition of unbroken (or exact) symmetry and emergent symmetry under an RG flow $\rho:\mathbf{A}_{(1)}\rightarrow\mathbf{A}_{(2)}$. A phenomenology of unbroken symmetry under RG flow is simple: a symmetry subring $\mathbf{A}_{(1)}^{\text{ub}}\subset\mathbf{A}_{(1)}$ of UV theory, which does not change under a RG flow $\rho$, is an unbroken symmetry. In the algebraic language, this condition means that $\mathbf{A}_{(1)}^{\text{ub}}$ and $\text{Ker}\rho$ are disjoint except for 0. Hence, the following defines the unbroken symmetry,
@@
$\begin{split}&\mathbf{A}_{(1)}^{\text{ub}}\text{ is an unbroken symmetry under }\rho\\ &\leftrightarrow\mathbf{A}_{(1)}^{\text{ub}}\cap\text{Ker}\rho=\{0\}\end{split}$ (10)
@@
In other words, when the above condition is not satisfied, the symmetry $\mathbf{A}_{(1)}^{\text{ub}}$ is deformed or broken nontrivially under $\rho$. This general argument provides a testable and accessible criterion of emergent and exact symmetry and will be useful to test conjectural RG flow, for example, in [123, 124, 125]. It should be remarked that as linear algebra, one can consider the basis of $\mathbf{A}_{(1)}/\text{Ker}\rho$, but it does not imply that $\mathbf{A}_{(1)}/\text{Ker}\rho$ forms subring of original ring $\mathbf{A}_{(1)}$. Hence, the dimension of unbroken symmetry (which should form a subalgebra) is less than the dimension of $\mathbf{A}_{(1)}/\text{Ker}\rho$ and the following inequality holds,
$\text{dim}\mathbf{A}_{(2)}\geq\text{dim}\mathbf{A}_{(1)}^{\text{ub}},$ (11)
@@
where dim is the dimension of a ring in linear algebra. By taking a maximal unbroken subalgebra under $\rho$, one can restrict the form of perturbation in the UV theory, and the above inequality provides the constraint in comparing the IR (or emergent) theory.
@@
For the readers interested in the perturbative quantum field theoretic description of our method, we note two possible research directions. The first is a straightforward analysis of the interpolation behavior of $H_{(1)}$ to $H_{(2)}$. This interpolation can be implemented by preparing the Hamiltonian $(1-\lambda)H_{(1)}+\lambda H_{(2)}$ with $\lambda=0$ to $\lambda=1$ and the perturbative analysis around $\lambda=0$ is useful also in QFTs. In this research direction, numerical methods such as the truncated conformal space approach[184, 185] and analysis of corresponding lattice models will be fundamentally important.
@@
The second is the detection of a maximal unbroken subalgebra in this subsection. By assuming that a general perturbation $H_{\text{per}}$ from a UV CFT (possibly, multiple perturbations in QFTs) is invariant under the symmetry corresponding to this subalgebra $\mathbf{A}_{(1)}^{\text{ub}}$, one can restrict the form of possible perturbations systematically. More conventionally, one can write down the condition as follows,
$H_{\text{per}},Q_{a^{\prime}_{(1)}}]=0,$ (12)
@@
where $Q_{a^{\prime}_{(1)}}$ is a symmetry operator (or integral of motions) corresponding to the element $a^{\prime}_{(1)}\in\mathbf{A}_{(1)}^{\text{ub}}$. Inversely, if the perturbation $H_{\text{per}}$ is given, one can determine possible RG flows by studying the maximal subalgebra structure which commutes with $H_{\text{per}}$ and its disjoint ideals. Related arguments based on group symmetry can be seen ubiquitously, and our argument can be considered as a generalization emphasizing the ideal structure. We note [186, 187] as references containing analysis of symmetry with ring structure and their action on bulk fields.
For the readers interested in higher-dimensional systems, we note that the above arguments themselves can be applied straightforwardly to $D$-dimensional systems when treating the generalized symmetries as integral of motion or symmetry operators. When the IR theory becomes an extended TQFT under massive perturbations, the subring structure formed by $d$-dimensional symmetry operators in UV theory, where d=0,,Dd=0,...,D specifies dimensionality of symmetry. This ring structure will be obtained from the application of the topological Wick rotation to composite defects in [188], and the subring symmetry of them will appear as symmetry of the IR theory. If the corresponding flows are massless, the quotient ring of UV theory appears as IR symmetry.
@@ II.2 Historical remarks on the projections in physics
Before moving to the next section, we comment on the history of the ideas to represent RG as a projection. The idea itself can be seen ubiquitously in the fields, but well-known earlier examples related to the discussions in this work are the corner-transfer method by Baxter [189](for more details, see the review[190]) and projective representation in Affleck-Kennedy-Lieb-Tasaki chain[46]. This kind of projection or projective representation played a significant role in studying many-body quantum systems both theoretically and numerically. Related historical aspects can be seen in the review [191].
Related to the argument in CFT in this manuscript, we also note the restriction of Hilbert space in quantum group symmetric models[192, 193, 194] which have a connection to the Dotsenko-Fateev Coulomb gas[195, 196] (or Feigin-Fuchs construction[197] based on Wakimoto free field representation[198]) and the massless RG flow[86]. More recently, the RG domain wall, which is in the main scope of this work, has been introduced in [31] and its general correspondence with existing coset CFTs has been proposed in [32]. Some of the properties of homomorphism, or the properties of symmetry-preserving domain wall, have already appeared in [32], and it is remarkable that this pioneering work appeared before the introduction of generalized symmetry[199]. Later, more general cases have been studied in [200, 201, 202, 203, 204] and more mathematical understanding has been developed in [120, 121, 115, 205, 67, 206, 207, 208]. The idea of RG domain wall has also been combined with the holography in [209, 210] and related bound of entropy of conformal interface has been discussed in [211]. More recently, the correspondence between symmetry-preserving domain wall[22, 23, 24, 25, 27] and RG domain wall has been realized in [206, 66, 87, 212, 68], but this is relatively less familiar in the fields.
@@ III Generating ideal: Role of nonabelian anyon or noninvertible symmetry
One can perform the exact construction by classifying ideals of a UV theory and obtain (infinitely) many new series of unexplored homomorphisms. In other words, if the above statement does not hold, the connection between the RG domain wall and massless flow will be broken. Further investigation of more detailed properties of the IR theory, such as conditions for the locality or nonlocality and conditions to obtain the IR theory with positive or integer fusion coefficients, is an open problem. However, surprisingly, we note that the related problems of the projection for Yang-Lee CFT flows have already been investigated much earlier in [225]. The categorical analog of the ideal called the additive congruence relation by ideal, will play a role in the more categorical description of our method [^2] We thank Jurgen Fuchs for teaching the terminology to us.. Related to this research direction, we remark that the idea of an ideal appears in the studies of Hopf algebra and related TQFTs (see [227, 228], for example). We also remark that in the studies on Hopf algebra, one can see the ideas to represent a RG as projection [229, 230, 231, 232] but their formalism is closer to Lagrangian formalism and its precise connection to RG domain wall is (reasonable but) unclear at this stage (see also the reviews[233, 234])
Figure 4: Picture of the noninvertible object (anyon or symmetry) as a seed to generate a condensable block of a theory. We used blue color for noninvertible objects, black color for fusion between the noninvertible object and other objects, and red color for objects in $\mathbf{A}_{1}$. Because of the noninvertible structure, one can send the ideal generated by $s_{(1)}$ to 0 only by applying a projection to the theory.
@@
In particular, if one takes a noninvertible object in $\mathbf{A}_{(1)}$, it generates a nontrivial ideal straightforwardly because the generated ideal cannot contain identity. From this view, one can say that a noninvertible symmetry corresponding to a nonabelian anyon is a seed to obtain an ideal and corresponding homomorphism or massless RG flow (FIG. 4). By combining this with Moore-Seiberg data[165, 166, 167, 168], the fundamental importance of further systematic studies on noninvertible symmetries (or integral of motions) and the corresponding nonabelian anyons for the general understanding of TOs becomes more evident.
One can construct possible IR theories only from the algebraic data of a UV theory, but the construction of the homomorphisms between given UV and IR theories requires more involved calculation and consistency checks. In the next section, we solve this problem in several models.
@@ IV Construction of homomorphism
In this section, we study ring homomorphisms between fusion rings appearing in massless RG flow of conformal field theories[86]. We study the massless RG flow between ultraviolet (UV) CFT $\mathbf{B}$ to infrared (IR) CFT $\mathbf{B}^{\prime}$ where $\mathbf{B}$ and $\mathbf{B}^{\prime}$ represent nonchiral fusion ring or anyon algebra of spherical fusion category (SFC)[235, 236, 157]. Roughly speaking, the SFC corresponds to the bosonic modular invariant and conformal bootstrap technique[237]. We concentrate our attention on the algebraic relation of objects or anyons, study the ring homomorphism $\rho$: $\mathbf{B}$ $\rightarrow$ $\mathbf{B}^{\prime}$ (or algebraic data of a functor between two SFCs[21, 23, 25]). To simplify the notation and discussion, we assume the theory is in the scope of Moore-Seiberg data[165, 166, 167, 168] and the ring homomorphism between nonchiral fusion rules naturally introduces a ring homomorphism between chiral fusion rules and vice versa. Algebraically, this is trivial because of the ring isomorphism between the nonchiral and chiral fusion rules.
Before moving into the respective discussions, we remark on one of the most nontrivial points of our findings. Surprisingly, even in the flow from the tricritical Ising model to the Ising model, the homomorphism is not unique. However, by requiring a variant of integrability, the solution becomes unique. This phenomenon itself is analogous to ideal number decomposition by Kummer, and the integrability will play a fundamental role in constructing a treatable class of models. Inversely, the other homomorphisms are mysterious to some extent, but there exists a conjectural phenomenology to them. If one assumes the IR theory $\mathbf{A}_{(2)}$ or consequence of RG flow is integrable, one can say this homomorphism induces partially integrable embedding of the theory to UV theory $\mathbf{A}_{(1)}$, $\mathbf{A}_{(2)}\subset\mathbf{A}_{(1)}$. Fortunately, there exist very similar phenomena called partially solvable systems in recent literature[91, 92, 93] [^3] We thank Chihiro Matsui for introducing us to the idea of this embedding and related ideas in the literature and Hosho Katsura for indicating useful references in this research field.. Historical aspects of those models have been summarized in the introduction of [93], and we note the Hilbert space fragmentation as a fundamental aspect, and note a recent review [239]. We conjecture that a ring homomorphism generally corresponds to integrable or partially solvable systems. Further analysis of the RG flows by perturbative QFT methods is an interesting direction, and it will be described by UV CFT with multiple perturbations as in [240].